You will be notified at trevant@gmail.com when a tutor answers your question. You may receive more than one answer.
Your question:
Use the elimination method to solve the system of equations.
a + 9b = 12
a + 2b = -2
Select the correct choice below and, if necessary, fill in the answer box to complete your choice.
A. There is one solution, it is ____
B. There are infinitely many solutions.
c. There is no solution

Answer by mananth(16946)

1 a + 9 b = 12 .............1
Total value
1 a + 2 b = -2 .............2
Eliminate a
(1)-(2)
9b-2b=12-(-2)
7b=14
/7
b=2
plug b in (2)
a+2b=-2
a+2*2=-2
a+4=-2
a=-6
It has only one solution